# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

### v2

@@ 3.4 Further improvements of color sums (“ihel3p1”)
4 Kernel splitting II: Feynman diagram groups
After implementing helicity-level parallelism, separating Feynman diagrams from color sums and porting the latter to BLAS, the next step in my kernel splitting developments consisted, quite naturally, in trying to further decompose the calculate_jamps kernel, which in ihel3 computes dual amplitudes from all Feynman diagrams, into many shorter kernels. My initial motivation for attempting this was to investigate whether this may lead to further throughput increases, possibly thanks to the lower number of registers used by smaller GPU kernels (lower “register pressure”): I soon realised, however, that I was not able to achieve further throughput increases, because moderately small Feynman diagram kernels are still quite complex and especially come at the cost of a very large increase in GPU global memory access, which massively slows down the calculation. However, there is also a second motivation to splitting Feynman diagrams, namely making the code easier to build and manage for very complex physics processes with very large numbers of diagrams. As described in this section, my work in this direction did achieve some interesting progress, such as the possibility to compute $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes which were previously non-computable in the CUDACPP software.
In the following subsections, I describe my second batch of three sets of kernel splitting changes, which I now also consider ready to be integrated in new production releases of CUDACPP: (ihel4) one kernel per Feynman diagram in each helicity stream; (ihel4p1) ditto, with optional CUDA graph orchestration; (ihel5) one kernel per diagram group in each helicity stream, with separate source code files; (ihel6) a single kernel in each helicity stream, processing all diagram groups as device functions. The results of the intermediate ihel4 step and those of the final ihel6p2 step on $gg\!\rightarrow\!t\bar{t}$ production with 0, 1, 2 or 3 extra final state gluons are given in Figures 4 and 5 for NVidia and AMD GPUs: these plots show, in particular, that the achieved throughputs in the ihel6p2 software are completely equivalent to those of the earlier ihel3p1 software, when the code is generated without diagram splitting, i.e. with a single kernel processing all Feynman diagrams as in ihel3p1. The main interest of ihel6p2, conversely, comes from the more complex $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$ with 4 and 5 extra final state gluons, the results for which are presented at the end of this section.
Figure 10: Throughputs (ME/s) as a function of grid size for an NVidia V100 GPU (node itscrd90 at CERN). Higher is better. The 2 plots correspond to throughputs derived by executing the check.exe application through 1 (left) or 10 (right) cycles of a GPU grid. The number of threads per block is fixed to 32 (warp size); the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel3) all Feynman diagrams in a single kernel; (ihel4) Feynman diagrams as individual kernels, without CUDA graphs; (ihel4p1) Feynman diagrams as individual kernels, with CUDA graphs.
4.1 Feynman diagrams as individual kernels (“ihel4”)
@@
There are at least two possible explanations for the worse GPU performance of ihel4: heavier access to GPU global memory, and kernel launch overhead. At the time of writing my previous preprint bib:preprintV1 for this article, I wrongly assumed that the latter was the most likely culprit. With hindsight and after many additional tests and developments, it is now clear that the bottleneck is instead the former. Kernel launch overhead can be reduced by adopting the CUDA graphs bib:graphsblog technology: this is what I implemented in the ihel4p1 software, described below. Reducing the impact of GPU global memory access, conversely, required a much more radical rethink of Feynman diagram splitting, described in the two following subsections: first of all, by launching kernels which include groups of many Feynman diagrams (ihel5), rather than individual diagrams; later on, by moving back to a single kernel, which however is no longer monolithic as it internally calls diagram groups as GPU device functions (ihel6).
Figure 11: Schematic representation of the CUDACPP ME engine (sigmaKin), in the last kernel splitting scenario described in this paper: (ihel6) Feynman diagram groups as individual kernels or as device functions in a single kernel.
@@ 4.1.1 Small kernels in CUDA graphs (“ihel4p1”)
The performance results for this implementation can be seen in Fig. 10, by comparing the grey (ihel4, no graphs), cyan (ihel4p1, graphs enabled) and red (ihel3, single kernel) curves, for a single execution (left) or ten executions (right) of the workflow. As expected, these curves indicate that graphs slow down the workflow when kernels are executed only once (ihel4p1 is worse than ihel4 in the left plots), but they speed it up when the workflow is executed several times (ihel4p1 is worse than ihel4 in the right plots). However, the plots also show that, even if graphs are used, launching Feynman diagrams as individual kernels is still much slower than keeping them together in a single kernel (ihel4p1 is much worse than ihel3 in both plots). This is a clear indication that the main problem in this approach is not the kernel launch overhead, but global memory access: this is precisely the issue addressed using diagram groups in the ihel5 and ihel6 versions of the software, which are described in the two next subsections.
@@ 4.2 Feynman diagram groups as kernels (“ihel5”)
After showing that the impact of kernel launch overhead was relatively limited, it was thus clear that the most likely explanation for the worse GPU performance of ihel4 with respect to ihel3 is memory access. One technical challenge in this case, in fact, is that the wavefunctions for all external and internal particles, which are a local variable wf in the calculate_wavefunctions and calculate_jamps kernels in versions ihel0–ihel3 of the code (see Table 1), must now also be stored into and retrieved from GPU global memory. In addition, as previously mentioned in a footnote, the existing implementation of helicity amplitude functions like FFV1_0 essentially constrains the wavefunction layout in GPU global memory to be an Array-Of-Structures (AOS) wf[nevt][6][2], which is suboptimal as it does not allow coalesced memory access as an SOA would do. This further increases the pressure on memory bandwidth, and may represent an additional cause of slowdowns.
Figure 12: Throughputs (ME/s) as a function of grid size for $gg\!\rightarrow\!t\bar{t}ggg$ on an NVidia V100 GPU. Higher is better. The number of threads per block is fixed to 32 (warp size); the grid size is varied by changing the number of blocks. Each plot compares code generated with different numbers of diagrams per group (dpg), ranging from 1 to 2000. The thick solid line (dpg2000) corresponds to the case where all 1240 diagrams are in a single diagram group, i.e. to the default scenario for $gg\!\rightarrow\!t\bar{t}ggg$ in the rest of this paper. The 2 plots correspond to code built in the DCDIAG=0 (left: each diagram group is a kernel) and DCDIAG=1 (right: there is a single kernel, calling all diagram groups as device functions) build modes.
@@
My ihel5 implementation of diagram groups is configurable according to two parameters: one “dpg” parameter defines the maximum number of Feynman diagrams in one diagram group, while another “dpf” parameter defines the minimum number of Feynman diagrams in one source code file. Currently these parameters are only managed through special scripts, but eventually they could become standard MG5aMC runcard parameters. In practice, diagram group kernels are implemented as follows: for each diagram group, the code generator keeps track of which wavefunctions the diagram group needs in its calculations, and of which wavefunctions the diagram group updates; at the beginning of all diagram groups except the first one, those wavefunctions are read from global memory and copied to local variables in GPU registers; incidentally, this makes it possible to define a better SOA layout in global memory for coalesced memory access, while maintaining the legacy AOS layout wf[nevt][6][2] in local variables to avoid the need to restructure all calculations; at the end of all diagram groups except the last one, any updated wavefunctions are written back to GPU global memory; at the end of all diagram groups, the jamps dual amplitudes in GPU global memory are also updated from the results of those computed using local variables. This design implies that, the lower the number of diagram groups, the lower the frequency of GPU global memory access. In particular, in the special case where the dpg is larger than the total number of diagrams in the physics process, only one diagram group is created: this is executed as a single kernel, which does not need to read wavefunctions at the beginning or to write them back at the end, so effectively this is a single-kernel implementation equivalent to that in ihel3, and indeed has the same performance, as previously noted.
For reasons clarified below, ihel5 is equivalent to the default DCDIAG=0 implementation in the ihel6 version of the software, which is illustrated in the upper-left part of Fig. 11. As a consequence, in this paper I do not show explicit results for the ihel5 software, but only those from ihel6. The impact of the dpg parameter on ihel5 performance for $gg\!\rightarrow\!t\bar{t}ggg$, in particular, is shown by the left plot in Fig. 12, derived from the default DCDIAG=0 version of ihel6. The plot is interesting as it shows that peak performance, equivalent to that of ihel3, can only be achieved with a single kernel that processes all diagrams: even a minimal splitting of the 1240 diagrams into two diagram groups (dpg=1000), in fact, comes at the cost of a very large performance hit for small GPU grids and of a 10% reduction in peak performance for large grids (with the additional problem that the program crashes due to limited memory when the GPU grids are too large). This clearly shows that GPU global memory access is very expensive.
@@ 4.3 Diagram groups as device functions (“ihel6”)
The impact on performance of the DCDIAG=1 approach is quite interesting, as shown in the right plot in Fig. 12. As in the case of DCDIAG=0, performance is quite poor for low values of dpg, i.e. when there are many diagram groups. Unlike DCDIAG=0, however, peak performance, only slightly lower than that of ihel3, is reached for intermediate values of dpg between 100 and 1000. For dpg=2000, instead, i.e. with a single diagram group, performance degrades again. In other words, there seems to be a sweet spot for diagram groups of intermediate sizes, where DCDIAG=1 performance is reasonably good. This is especially interesting because this is the same sweet spot where source code files and their builds appear to be easier to manage. This is further discussed in the following, in the case of the more complex processes $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$.
@@
Table 4: C++ run time performance of $gg\!\rightarrow\!t\bar{t}gggg$ in different SIMD modes and with different numbers of diagrams per group (FPTYPE=m). C++ code generated using CUDACPP hack_ihel6p2. Results for a single core of an Intel Xeon Gold 6326 CPU. The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering, which is still performed sequentially one helicity at a time. As a basic validation of physics results, it was only checked that all calculations return exactly the same average ME value, when starting from the same random number seeds, for all SIMD and diagram-per-group modes.
@@
Table 5: CUDA run time performance of $gg\!\rightarrow\!t\bar{t}gggg$ in different DCDIAG modes and with different numbers of diagrams per group (FPTYPE=m). CUDA/C++ code generated using CUDACPP ihel6p2. Results for an NVidia A100 GPU. The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering, which is still performed sequentially one helicity at a time. As a basic validation of physics results, it was only checked that all calculations return exactly the same average ME value, when starting from the same random number seeds, for all SIMD and diagram-per-group modes.
@@
Before moving on to more complex physics processes that can now be computed by splitting the calculation into several diagram groups, I add one final comment about our standard candles, $g\!\rightarrow\!t\bar{t}, gg$$\rightarrow\!t\bar{t}g, gg$$g\!\rightarrow\!t\bar{t}gg$ and $gg\!\rightarrow\!t\bar{t}ggg$. In particular, as shown in Figures 4 and 5 for NVidia and AMD GPUs and in Fig. 6 for an Intel CPU, I stress that the ihel6 line of software versions, and more specifically the “ihel6p2” version described below, performs just as well as the current ihel3p1 release, or even slightly better, when these processes are generated using a single diagram group. In other words, the new diagram splitting functionality in ihel6p2 offers the benefit of allowing computations of more complex processes, without compromising the well established performance for simpler processes.
4.4 Studies of a $2\!\rightarrow\!6$ process: $gg\!\rightarrow\!t\bar{t}gggg$
Using the new features in CUDACPP ihel6, notably its diagram splitting capabilities and the possibility to treat diagram groups either as independent kernels or as device methods on GPUs, I was able to test and obtain some throughput results for complex $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ physics processes such as $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$. I note that some studies of $gg\!\rightarrow\!t\bar{t}gggg$ had already been performed in the past using the previous ihel0 software with a monolithic sigmaKin kernel, by myself and other members of the development team, but they had immediately appeared extremely challenging. In my recollection, in 2023 I had been able to build the C++ code for $gg\!\rightarrow\!t\bar{t}gggg$, but only using the clang compiler, and this had taken 32 hours. The C++ build using g++, conversely, had crashed with an internal error, while the CUDA build had gone on for more than one week before I decided to stop it.
Diagram splitting dramatically improves this situation, not only for the GPU implementation that I was initially targeting, but also for the vectorised C++ implementation. As shown in Table 4, the C++ build of $gg\!\rightarrow\!t\bar{t}gggg$ with g++ successfully completes in less than 4 minutes if the number dpg of diagrams per group is kept below 1000, and it takes longer but still completes in 30 minutes if dpg is increased to 10000, which implies only two diagram groups for this specific process. Interestingly, instead, the g++ build fails with an internal compiler error if all 15495 diagrams are included in the same function: this is exactly the same situation that I had faced earlier on with the ihel0 software. Table 4 is also very interesting because it shows that C++ runtime throughputs are significantly degraded with fewer than 100 dpg, but they are essentially at the their peak plateau with 1000 dpg: the runtime performance with 10k dpg is only negligibly better than that with 1k dpg, at the cost of a much longer build time. This observation, which seems to fall in line with that previously made about $gg\!\rightarrow\!t\bar{t}ggg$ for GPUs, is especially interesting because it seems to suggest that, if the C++ build of the 100k dpg code (i.e. that with a monolithic function) had succeeded, its runtime performance might have been only negligibly better than that with 10k or 1k dpg. In other words, also for the C++ implementation of $gg\!\rightarrow\!t\bar{t}gggg$ (similarly to what was previously observed for the GPU implementation of $gg\!\rightarrow\!t\bar{t}ggg$), there seems to be a sweet spot around 1k dpg where software builds are fast and manageable, and runtime performance is satisfactory. It should be noted, in particular, that the SIMD speedup observed with dpg 1k or 10k is consistent with the maximum achievable in this configuration, using mixed floating point precision: on this Intel Xeon Gold 6326 CPU with two FMA units, one would expect a maximum SIMD speedup of x8 for Feynman diagrams and x16 for the color matrix, and an overall speedup somewhere in between, which is consistent with the observed speedups around x9. This very good scaling seems to suggest that this calculation is not limited by memory access on the CPU.
@@ 4.4 Studies of a $2\!\rightarrow\!6$ process: $gg\!\rightarrow\!t\bar{t}gggg$
Table 6: Profiling of Feynman diagrams and color sums for $gg\!\rightarrow\!t\bar{t}gggg$ on CPU in different SIMD modes and with different numbers of diagrams per group (FPTYPE=m). C++ code generated using a modified version of CUDACPP, instrumented with rdtsc timers (hack_ihel6p2_ggtt4g branch). Results for a single core of an Intel Xeon Gold 6326 CPU. The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering. The average ME value shown in the table, computed from the same random numbers in all cases, is used as a basic validation of physics results.

### v3

@@ 4.1 Feynman diagrams as individual kernels (“ihel4”)
There are at least two possible explanations for the worse GPU performance of ihel4: heavier access to GPU global memory, and kernel launch overhead. At the time of writing my first preprint bib:preprintV1 for this article, I wrongly assumed that the latter was the most likely culprit. With hindsight and after many additional tests and developments, it is now clear that the bottleneck is instead the former. Kernel launch overhead can be reduced by adopting the CUDA graphs bib:graphsblog technology: this is what I implemented in the ihel4p1 software, described below. Reducing the impact of GPU global memory access, conversely, required a much more radical rethink of Feynman diagram splitting, described in the two following subsections: first of all, by launching kernels which include groups of many Feynman diagrams (ihel5), rather than individual diagrams; later on, by moving back to a single kernel, which however is no longer monolithic as it internally calls diagram groups as GPU device functions (ihel6).
Figure 10: Throughputs (ME/s) as a function of grid size for an NVidia V100 GPU (using CUDA 12.0 and gcc 11.5). Higher is better. The 2 plots correspond to throughputs derived by executing the check.exe application through 1 (left) or 10 (right) cycles of a GPU grid. The number of threads per block is fixed to 32 (warp size); the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel3) all Feynman diagrams in a single kernel; (ihel4) Feynman diagrams as individual kernels, without CUDA graphs; (ihel4p1) Feynman diagrams as individual kernels, with CUDA graphs.
@@ 4.1.1 Small kernels in CUDA graphs (“ihel4p1”)
The performance results for this implementation are shown in Fig. 10, which compares three grey (ihel4, no graphs), cyan (ihel4p1, graphs enabled) and red (ihel3, single kernel) curves, for a single execution (left) or ten executions (right) of the workflow. As expected, these plots indicate that graphs slow down the workflow when kernels are executed only once (ihel4p1 is worse than ihel4 in the left plots), but they speed it up when the workflow is executed several times (ihel4p1 is worse than ihel4 in the right plots). However, Fig. 10 also shows that, even if graphs are used, launching Feynman diagrams as individual kernels is still much slower than keeping them together in a single kernel (ihel4p1 is much worse than ihel3 in both plots). This is a clear indication that the main problem in this approach is not the kernel launch overhead, but global memory access: this is precisely the issue addressed using diagram groups in the ihel5 and ihel6 versions of the software, which are described in the two next subsections.
Figure 11: Schematic representation of the CUDACPP ME engine (sigmaKin), in the last two kernel splitting scenarios described in this paper: (ihel5, and ihel6 with DCDIAG=0) Feynman diagram groups as individual kernels; (ihel6 with DCDIAG=1) Feynman diagram groups as device functions in a single kernel. The representation of color sums is valid for all versions from ihel3p1 onwards.
@@ 4.2 Feynman diagram groups as kernels (“ihel5”)
After showing that the impact of kernel launch overhead is relatively limited, it is thus clear that the most likely explanation for the worse GPU performance of ihel4 with respect to ihel3 is memory access. One technical challenge in this case, in fact, is that the wavefunctions for all external and internal particles, which are a local variable wf in the calculate_wavefunctions and calculate_jamps kernels in versions ihel0–ihel3 of the code (see Table 1), must now also be stored into and retrieved from GPU global memory. In addition, as previously mentioned in a footnote, the existing implementation of helicity amplitude functions like FFV1_0 essentially constrains the wavefunction layout in GPU global memory to be an Array-Of-Structures (AOS) wf[nevt][6][2], which is suboptimal as it does not allow coalesced memory access as an SOA would do. This further increases the pressure on memory bandwidth, and may represent an additional cause of slowdowns.
Figure 12: Throughputs (ME/s) as a function of grid size for $gg\!\rightarrow\!t\bar{t}ggg$ on an NVidia V100 GPU (using CUDA 12.0 and gcc 11.5). Higher is better. The number of threads per block is fixed to 32 (NVidia GPU warp size); the grid size is varied by changing the number of blocks. Each plot compares code generated with different numbers of diagrams per group (dpg), ranging from 1 to 2000. The thick solid line (dpg2000) corresponds to the case where all 1240 diagrams are in a single diagram group, i.e. to the default scenario for $gg\!\rightarrow\!t\bar{t}ggg$ elsewhere in this paper. The 2 plots correspond to code built in the DCDIAG=0 (left: each diagram group is a kernel) and DCDIAG=1 (right: there is a single kernel, calling all diagram groups as device functions) build modes.
@@
My ihel5 implementation of diagram groups is configurable according to two parameters: one “dpg” parameter defines the maximum number of Feynman diagrams in one diagram group, while another “dpf” parameter defines the minimum number of Feynman diagrams in one source code file. Currently these parameters are only managed through special scripts, but eventually they could become standard MG5aMC runcard parameters. In practice, diagram group kernels are implemented as follows: for each diagram group, the code generator keeps track of which wavefunctions the diagram group needs in its calculations, and of which wavefunctions the diagram group updates; at the beginning of all diagram groups except the first one, all necessary wavefunctions are read from global memory and copied to local variables in GPU registers; incidentally, this makes it possible to define a better SOA layout in global memory for coalesced memory access, while maintaining the legacy AOS layout wf[nevt][6][2] in local variables to avoid the need to restructure all calculations; at the end of all diagram groups except the last one, any updated wavefunctions are written back to GPU global memory; at the end of all diagram groups, the jamps dual amplitudes in GPU global memory are also updated from the results of those computed using local variables. This design implies that, the lower the number of diagram groups, the lower the frequency of GPU global memory access. In particular, in the special case where the dpg is larger than the total number of diagrams in the physics process, only one diagram group is created: this is executed as a single kernel, which does not need to read wavefunctions at the beginning or to write them back at the end, so effectively this is a single-kernel implementation equivalent to that in ihel3, and indeed has the same performance, as previously noted.
For reasons clarified below, ihel5 is equivalent to the default DCDIAG=0 implementation in the ihel6 version of the software, which is illustrated in the upper-left part of Fig. 11. As a consequence, in this paper I do not show explicit results for the ihel5 software, but only those from ihel6. The impact of the dpg parameter on ihel5 performance for $gg\!\rightarrow\!t\bar{t}ggg$, in particular, is shown by the left plot in Fig. 12, derived from the default DCDIAG=0 version of ihel6. The plot is interesting as it shows that peak performance, equivalent to that of ihel3, can only be achieved with a single kernel that processes all diagrams. Even a minimal splitting of the 1240 diagrams into two diagram groups (dpg=1000), in fact, comes at the cost of a very large performance hit for small GPU grids and of a 10% reduction in peak performance for large grids (with the additional problem that the program crashes due to limited memory when the GPU grids are too large). This clearly shows that GPU global memory access is very expensive.
@@ 4.3 Diagram groups as device functions (“ihel6”)
The impact on performance of the DCDIAG=1 approach is quite interesting, as shown in the right plot in Fig. 12. As in the case of DCDIAG=0, performance is quite poor for low values of dpg, i.e. when there are many diagram groups. Unlike DCDIAG=0, however, peak performance, only slightly lower than that of ihel3, is reached for intermediate values of dpg between 100 and 1000. For dpg=2000, instead, i.e. with a single diagram group, performance degrades again. In other words, there seems to be a sweet spot for diagram groups of intermediate sizes, where DCDIAG=1 performance is reasonably good. This is especially interesting because it is the same sweet spot where source code files and their builds appear to be easier to manage. This is further discussed in the following, in the case of $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$, which can now be computed by splitting the calculation into several diagram groups.
@@
Table 5: C++ run time performance of $gg\!\rightarrow\!t\bar{t}gggg$ in different SIMD modes and with different numbers of diagrams per group (FPTYPE=m). C++ code generated using CUDACPP hack_ihel6p2. Results for a single core of an Intel Xeon Gold 6326 CPU (using gcc 11.5). The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering, which is still performed sequentially one helicity at a time. As a basic validation of physics results, it was only checked that all calculations return exactly the same average ME value, when starting from the same random number seeds, for all SIMD and diagram-per-group modes.
@@
Table 6: CUDA run time performance of $gg\!\rightarrow\!t\bar{t}gggg$ in different DCDIAG modes and with different numbers of diagrams per group (FPTYPE=m). CUDA/C++ code generated using CUDACPP ihel6p2. Results for an NVidia A100 GPU (using CUDA 12.4 and gcc 11.5). The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering, which is still performed sequentially one helicity at a time. As a basic validation of physics results, it was only checked that all calculations return exactly the same average ME value, when starting from the same random number seeds, for all DCDIAG and diagram-per-group modes.
@@
Before moving on to these more complex physics processes, I add one final comment about our standard candles, $g\!\rightarrow\!t\bar{t}, gg$$\rightarrow\!t\bar{t}g, gg$$g\!\rightarrow\!t\bar{t}gg$ and $gg\!\rightarrow\!t\bar{t}ggg$. As shown in Figures 4 and 5 for NVidia and AMD GPUs and in Fig. 6 for an Intel CPU, what I want to stress is that the ihel6 line of software versions, and more specifically the “ihel6p2” version described below, performs just as well as the current ihel3p1 release, or even slightly better, when these processes are generated using a single diagram group. In other words, the new diagram splitting functionality in ihel6p2 offers the benefit of allowing computations of more complex processes, without compromising the well established performance for simpler processes.
@@
Table 7: Profiling of Feynman diagrams and color sums for $gg\!\rightarrow\!t\bar{t}gggg$ on CPU in different SIMD modes and with different numbers of diagrams per group (FPTYPE=m). C++ code generated using a modified version of CUDACPP, instrumented with rdtsc timers (hack_ihel6p2_ggtt4g branch). Results for a single core of an Intel Xeon Gold 6326 CPU (using gcc 11.5). The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering. The average ME value shown in the table, computed from the same random numbers in all cases, is used as a basic validation of physics results.
4.4 Studies of a $2\!\rightarrow\!6$ process: $gg\!\rightarrow\!t\bar{t}gggg$
Using the new features in CUDACPP ihel6, notably its diagram splitting capabilities and the possibility to treat diagram groups either as independent kernels or as device methods on GPUs, I was able to test and obtain some throughput results for complex $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ physics processes such as $gg\!\rightarrow\!t\bar{t}gggg$ and $gg\!\rightarrow\!t\bar{t}ggggg$. I note that some studies of $gg\!\rightarrow\!t\bar{t}gggg$ had already been performed in the past using the previous ihel0 software with a monolithic sigmaKin kernel, by myself and other members of the development team, but they had immediately appeared extremely challenging. In my recollection, in 2023 I had been able to build the C++ code for $gg\!\rightarrow\!t\bar{t}gggg$, but only using the clang compiler, and this had taken 32 hours. The C++ build using gcc, conversely, had crashed with an internal error, while the CUDA build had gone on for more than one week before I decided to stop it.
Diagram splitting dramatically improves this situation, not only for the GPU implementation that I was initially targeting, but also for the vectorized C++ implementation. As shown in Table 5, the C++ build of $gg\!\rightarrow\!t\bar{t}gggg$ with gcc successfully completes in less than 4 minutes if the number dpg of diagrams per group is kept below 1000, and it takes longer but still completes in 30 minutes if dpg is increased to 10000, which implies only two diagram groups for this specific process. Interestingly, instead, the gcc build fails with an internal compiler error if all 15495 diagrams are included in the same function: this is exactly the same situation that I had faced earlier on with the ihel0 software. Table 5 is also very interesting because it shows that C++ runtime throughputs are significantly degraded with fewer than 100 dpg, but they are essentially at the their peak plateau with 1000 dpg: the runtime performance with 10k dpg is only negligibly better than that with 1k dpg, at the cost of a much longer build time. This observation, which seems to fall in line with that previously made about $gg\!\rightarrow\!t\bar{t}ggg$ for GPUs, is especially interesting because it seems to suggest that, if the C++ build of the 100k dpg code (i.e. that with a monolithic function) had succeeded, its runtime performance might have been only negligibly better than that with 10k or 1k dpg. In other words, also for the C++ implementation of $gg\!\rightarrow\!t\bar{t}gggg$ (similarly to what was previously observed for the GPU implementation of $gg\!\rightarrow\!t\bar{t}ggg$), there seems to be a sweet spot around 1k dpg where software builds are fast and manageable, and runtime performance is satisfactory. It should be noted, in particular, that the SIMD speedup observed with dpg 1k or 10k is consistent with the maximum achievable in this configuration, using mixed floating point precision: on this Intel Xeon Gold 6326 CPU with two FMA units, one would expect a maximum SIMD speedup of x8 for Feynman diagrams and x16 for the color matrix, and an overall speedup somewhere in between, which is consistent with the observed speedups around x9. This very good scaling seems to suggest that this calculation is not limited by memory access on the CPU.